OpenStudy (anonymous):

For the polynomial function, find the specified function values. f(x)=3x^2+3x^2-2 find f(0) and f(-1).

OpenStudy (mertsj):

f(0)=-2 f(-1)=4

OpenStudy (anonymous):

f(x) = 3x^2 + 3x^2 - 2 f(0) = 3(0)^2 + 3(0)^2 - 2 f(0) = 0 + 0 - 2 f(0) = -2 f(x) = 3x^2 + 3x^2 - 2 f(-1) = 3(-1)^2 + 3(-1)^2 - 2 f(-1) = 3 + 3 - 2 f(-1) = 4